OpenStudy (anonymous):

find value of x to make statement true 3/5x - 5+x-4=-1

OpenStudy (anonymous):

^ if that is a fraction. thats what I assumed

OpenStudy (mathstudent55):

@BVB4rmy Almost. Here are your first two lines: "3/5x-5+x-4=-1 3/5x+x=9" Let's do it one step at a time: 3/5x-5+x-4=-1 3/5x + x - 5 - 4 = -1 3/5x + x - 9 = -1 3/5x + x = -1 + 9 3/5x + x = 8 The line above is what your second line should read.

OpenStudy (anonymous):

الآن فهمتك وأنت محق o.o

OpenStudy (anonymous):

:3 true

OpenStudy (mathstudent55):

Of course, we're both assuming the equation is: \(\dfrac{3}{5}x - 5 + x - 4 = -1\)